# Demographic-aware fine-grained classification of pediatric wrist fractures

### v4

@@ Abstract
Wrist pathologies are frequently observed, particularly among children who constitute the majority of fracture cases. Computer vision presents a promising avenue, contingent upon the availability of extensive datasets, a notable challenge in medical imaging. Therefore, reliance solely on one modality, such as images, proves inadequate, especially in an era of diverse and plentiful data types. This study addresses the problem using a multifaceted approach: framing it as a fine-grained recognition task, fusing patient metadata with X-rays, and leveraging weights from a separate fine-grained dataset rather than from a coarse-grained dataset like ImageNet. Unlike prior work, this is the first application of metadata integration for wrist pathology recognition. Our results show that combining fine-grained transformer approach, fine-grained pre-training, and metadata integration improves diagnostic accuracy by 2% on small custom curated dataset and over 10% on a larger fracture dataset.
@@ 1 Introduction
Hand injuries are among the most frequent presentations in emergency departments, with wrist fractures accounting for up to 75% of cases [1]. Given the wrist’s complex anatomy and critical role in hand function, accurate diagnosis is essential to prevent long-term impairment. Recent advances in computer vision have shown promise in detecting trauma-related abnormalities on X-rays; however, such methods typically require large, well-annotated datasets, which remain scarce in medical imaging. To overcome this limitation, multimodal learning approaches that combine patient metadata with imaging have been explored in domains such as dermatology and oncology, showing consistent improvements in diagnostic accuracy [2, 3, 4, 5, 6].
A consistent trend in these studies is the predictive value of demographic metadata, particularly age and sex. Thomas et al. [6] demonstrated accuracy gains by fusing age, sex, and anatomical site with dermoscopy images, while Li et al. [5] reported 89.1% accuracy on ISIC 2018 using PNASNet with metadata fusion. Pacheco et al. [7, 8] showed balanced accuracy improvements and introduced specialized fusion modules, and Ningrum et al. [2] achieved 92.3% melanoma classification by coupling ANN-based metadata encoders with CNNs. More recently, Cai et al. [4] applied ViTs with metadata encoders, reaching 93.8% accuracy. Collectively, these works establish metadata as a critical complement to imaging in medical AI.
Despite this progress, metadata integration has not been investigated for wrist pathology classification. This omission is critical because wrist anatomy and pathology presentation are strongly age- and sex-dependent. Carpal ossification occurs in well-defined stages, the distal radial and ulnar physes close at different ages for boys and girls [9, 10], and common pediatric fracture types (e.g., buckle, greenstick) differ from adult patterns [11, 12]. Without age/sex context, a model might misclassify normal developmental features (like unfused growth plates) as abnormalities.
To address this gap, we extend metadata integration to musculoskeletal imaging. We first frame wrist abnormality recognition as a fine-grained visual recognition (FGVR) task, unlike prior multimodal efforts that primarily used conventional CNNs. On top of this, we fuse demographic metadata with radiographs to capture subtle, age- and sex-specific morphological cues that image-only models often overlook. Building on our prior work image-only fine-grained recognition work [13, 14], we propose a joint framework for pediatric wrist pathology analysis and hypothesize that it will outperform image-only and CNN-based baselines.
@@ 2.1 Dataset Curation & Patient Metadata
We curated two datasets from GRAZPEDWRI-DX [15]. (1) A “limited” set of three wrist pathologies (fracture, bone anomaly, soft tissue), excluding the “metal” class, following our prior work [13, 14]. (2) A binary “Fracture vs. No Fracture” dataset using all images, with the non-fracture class augmented for balance. Final splits are shown in Table 1. The cohort included 6,091 patients (2,688 female, 3,402 male, ages 0.2–19 years, mean 10.9). Fig. 1 shows distributions: fractures are more common in males, while anomalies concentrate in older groups.
@@ 2.2 Architectural Details
This work utilizes a hybrid architecture that integrates both visual data and meta-information, drawing inspiration from the architecture called “MetaFormer” [16]. As depicted in Fig. 2, the architecture is organized into five stages. The first stage (S0) starts with a basic 3-layer convolutional stem, which processes raw input images. Subsequent stages progressively downsample the feature maps and increase channel dimensionality. Stages S1 and S2 consist of MBConv blocks enhanced with squeeze-and-excitation mechanism. The final two stages, S3 and S4, transition to transformer-based blocks that employ relative positional encoding. This design allows the model to capture global dependencies, which are crucial for subtle pathological differences that span larger contexts (e.g., comparing radius vs. ulna contours or detecting diffuse soft tissue swelling).
@@
Table 2: Configuration of the architecture series. B represents the number of blocks, and H denotes the hidden dimensions for each stage.
@@
The final output $y$ aggregates information across multiple scales and token types. When using external attributes as metadata, these are encoded as vectors and embedded using a fully connected non-linear function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$. In our setup, the demographic attributes consist of age and gender. The age attribute was normalized using min–max scaling, while the gender attribute was one-hot encoded. This results in a 2D input vector, which is then projected into the same embedding space as the vision tokens, ensuring consistency between visual and meta representations.
@@ 2.3 Experimental Settings
A range of well-established models has been incorporated in our study alongside our fine-grained metadata-aware approach for comparison. All these baseline models were pre-trained on ImageNet [19]. These include ViT [20], VGG16 [21], GoogleNet [22], EfficientNetV2 [23], AlexNet [24], DenseNet201 [25], InceptionV3 [26], and ResNet50 [27]. All models underwent training for 100 epochs using an image resolution of 224x224 and a batch size of 32. Adam optimizer was utilized with a learning rate $\eta=5\times 10^{-3}$. For MetaFormer, however, the base learning rate was $\eta=1\times 10^{-4}$, warmup learning rate $\alpha=5\times 10^{-7}$, minimum learning rate was $\eta_{m}=1\times 10^{-5}$, and weight decay was $\wp=5\times 10^{-2}$.
@@ 3 Results & Discussion
ViT 62.71%
VGG16 63.28%
GoogleNet 70.34%
EfficientNetV2 72.32%
AlexNet 74.29%
DenseNet201 76.84%
InceptionV3 77.44%
ResNet50 77.68%
MetaFormer FGVR (Base) 79.40%
Table 4: Accuracy of different MetaFormer configurations on only image data, integrating metadata through early fusion, and integrating metadata through late fusion.
Model Configuration Vision Only Early Fusion Late Fusion
@@ 3.1 Baseline Comparison
Table 3 shows that MetaFormer (79.4%) outperforms widely used CNNs such as ResNet50 (77.7%) and DenseNet201 (76.8%). This gives credence to the fact that transformer-based architectures with fine-grained design are better suited to capture subtle morphological cues in pediatric wrist radiographs, where differences between normal anatomy and pathology can be minimal.
@@ 3.2 Metadata Integration & Pretraining
We then selected the “FG-2” configuration and used the weights obtained from training it on the iNaturalist dataset [28] to create our “FG-2-inat” configuration. This surpassed all other configurations, whether using image data alone or in conjunction with metadata. This configuration reached 81.4%, a 1.5% improvement over image-only training. It seems that the fine-grained pretraining improved generalization beyond ImageNet pretraining, reflecting the value of transferring from domains where subtle inter-class variations dominate. Table 5 shows per class sensitivity, specificity, and precision scores for configuration “FG-2-inat”.
Fig. 3: Heatmaps generated from the final convolutional layer in stage 2 (S2) of our configuration “FG-2-inat” on a limited dataset.
3.3 Interpretability
Table 6: Table showing accuracies obtained by our configuration of MetaFormer on individual metadata attributes and when both attributes are considered.
@@ 4 Conclusion
In this study, we demonstrated that integrating patient demographic metadata with image data enhances fine-grained recognition of pediatric wrist pathologies. Overall, our experiments demonstrate three key findings: (1) Fine-grained transformer architectures outperform conventional CNNs for pediatric wrist pathology. (2) Demographic metadata (age, sex) significantly improves recognition, with late fusion being most effective. (3) Fine-grained pretraining enhances generalization, and metadata boosts performance across both limited and large-scale datasets. These results emphasize that integrating developmental context into vision models is not only beneficial but essential for robust pediatric musculoskeletal diagnosis.
@@ 5 Acknowledgement
This work was supported by the Curricula Development and Capacity Building in Applied Computer Science for Pakistani Higher Education Institutions (CONNECT) Project NORPART-2021/10502, funded by the Norwegian Directorate for Higher Education and Skills (DIKU).

### v5

@@ Abstract
Wrist pathologies are frequently observed, particularly among children who constitute the majority of fracture cases. Computer vision presents a promising avenue, contingent upon the availability of extensive datasets, a notable challenge in medical imaging. Therefore, reliance solely on one modality, such as images, proves inadequate. This study addresses the problem using a multifaceted approach: framing it as a fine-grained recognition task, fusing patient metadata with X-rays, and leveraging weights from a separate fine-grained dataset rather than from a coarse-grained dataset like ImageNet. Unlike prior work, this is the first application of metadata integration for wrist pathology recognition. Our results show that our approach improves diagnostic accuracy by 2% on small custom curated dataset and over 10% on a larger fracture dataset.
@@ 1 Introduction
Wrist fractures represent up to 75% of hand injuries in emergency departments and pose significant risk of long-term impairment if misdiagnosed [1]. Given the wrist’s complex anatomy and developmental variability, accurate radiographic interpretation is essential. Recent advances in computer vision have demonstrated promise for automated abnormality detection in musculoskeletal imaging. However, such methods typically rely on large, labeled datasets, which are scarce in medical domains, and often process X-rays without contextual patient information.
To overcome the limitations of image-only learning, multimodal frameworks that fuse patient metadata with imaging have been successfully applied in dermatology and oncology. Studies have shown consistent accuracy improvements by incorporating age, sex, and anatomical site information alongside dermoscopic or histopathology images [2, 3, 4, 5, 6]. For example, Thomas et al. [6] fused demographic variables with dermoscopy images, while Cai et al. [4] applied ViTs with metadata encoders, achieving 93.8% melanoma classification accuracy. Recently, Hinterwimmer et al. [7] extended this approach to musculoskeletal imaging, reporting gains in bone tumour classification by integrating X-rays and metadata.
Despite these advances, metadata integration has not been investigated for wrist pathology recognition. This gap is significant, as age- and sex-specific skeletal variations critically influence wrist imaging. Carpal ossification follows predictable stages, growth plate closure occurs at different ages across sexes [8, 9], and pediatric fracture types differ from adult patterns [10]. Without demographic context, models may misclassify normal developmental features, such as unfused growth plates, as abnormalities.
We address this gap by (1) introducing the first framework that fuses radiographs with demographic metadata, (2) formulating the task as fine-grained recognition using transformer-based modeling, and (3) showing that fine-grained pretraining further boosts generalization rather than pretraining on ImageNet.
@@ 2.1 Dataset Curation & Patient Metadata
We curated two datasets from GRAZPEDWRI-DX [11, 12]. (1) A “limited” set of three wrist pathologies (fracture, bone anomaly, soft tissue), excluding the “metal” class, following our prior work [13]. (2) A binary “Fracture vs. No Fracture” dataset using all images, with the non-fracture class augmented for balance. Final splits are shown in Table 1. The cohort included 6,091 patients (2,688 female, 3,402 male, ages 0.2–19 years, mean 10.9). Fig. 2 shows distributions: fractures are more common in males, while anomalies concentrate in older groups.
@@ 2.2 Architectural Details
This work utilizes a hybrid architecture that integrates both visual data and meta-information, drawing inspiration from the MetaFormer architecture [14]. As depicted in Fig. 1, the architecture is organized into five stages. The first stage (S0) starts with a basic 3-layer convolutional stem, which processes raw input images. Subsequent stages progressively downsample the feature maps and increase channel dimensionality. Stages S1 and S2 consist of MBConv blocks enhanced with squeeze-and-excitation mechanism. The final two stages, S3 and S4, transition to transformer-based blocks that employ relative positional encoding. This design allows the model to capture global dependencies, which are crucial for subtle pathological differences that span larger contexts (e.g., comparing radius vs. ulna contours or detecting diffuse soft tissue swelling).
@@
Table 2: Configuration of the architecture series. B: number of blocks, and H: hidden dimensions for each stage.
@@
The final output $y$ aggregates information across multiple scales and token types. When using external attributes as metadata, these are encoded as vectors and embedded using a fully connected non-linear function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$.
@@ 2.3 Experimental Settings
We compare against common CNNs pre-trained on ImageNet [17, 18, 19, 20, 21, 22, 23, 24, 25], including ViT, VGG16, GoogleNet, EfficientNetV2, AlexNet, DenseNet201, InceptionV3, and ResNet50. All models underwent training for 100 epochs using a resolution of 224x224 and a batch size of 32. Adam optimizer was utilized with a learning rate $10^{-3}$. For MetaFormer, however, the base learning rate was $10^{-4}$.
@@ 3 Results & Discussion
ViT 62.7%
VGG16 63.3%
GoogleNet 70.3%
EfficientNetV2 72.3%
AlexNet 74.3%
DenseNet201 76.8%
InceptionV3 77.4%
ResNet50 77.7%
MetaFormer FGVR (Base) 79.4%
Table 4: Accuracy of different MetaFormer configurations on vision, vision + meta through early fusion, and late fusion.
Config Vision Early Fusion Late Fusion
@@ 3.1 Baseline Comparison
Table 3 shows that MetaFormer outperforms widely CNNs. This gives credence to the fact that transformer-based architectures with fine-grained design are better suited to capture subtle morphological cues in pediatric wrist radiographs, where differences between normal anatomy and pathology can be minimal.
@@ 3.2 Metadata Integration & Pretraining
Fig. 3: Heatmaps from the final convolutional layer in stage 2 (S2) of our configuration “FG-2-inat” on a limited dataset.
Table 6: Accuracies obtained by our configuration of MetaFormer on individual metadata attributes and when both attributes are considered.
@@
We then selected the “FG-2” configuration and used the weights obtained from training it on the iNaturalist dataset [26] to create our “FG-2-inat” configuration. This surpassed all other configurations, whether using image data alone or in conjunction with metadata. This configuration reached 81.4%, a 1.5% improvement over image-only training. It seems that the fine-grained pretraining improved generalization beyond ImageNet pretraining, reflecting the value of transferring from domains where subtle inter-class variations dominate. Table 5 shows per-class sensitivity, specificity, & precision scores for configuration “FG-2-inat”. Similarly, as shown in Fig. 4, metadata enhances fracture detection while preserving performance in bone anomaly and soft tissue recognition.
Fig. 4: ROC curves showing that metadata substantially improves fracture discrimination (AUC 0.84→0.96).
3.3 Interpretability
@@ 4 Conclusion
Overall, our experiments demonstrate three key findings: (1) Fine-grained transformer architecture outperform conventional CNNs for pediatric wrist pathology. (2) Demographic metadata (age, sex) significantly improves recognition across both limited and large-scale datasets, with late fusion being most effective. (3) Fine-grained pretraining enhances generalization compared to using ImageNet weights. These results emphasize that integrating developmental context into vision models is essential for robust pediatric musculoskeletal diagnosis.
@@ 5 Acknowledgement
This work was supported by the CONNECT Project NORPART-2021/10502, funded by the Norwegian Directorate for Higher Education and Skills (DIKU).